Give the slope of the line that passes through the points (-9,7) and (-2,-8)
Answer by nerdybill(7384)

For any two points (on a straight line)
(x1,y1) and (x2,y2)
the slope is
m = (y2-y1)/(x2-x1)
.
In your problem, the two points were:
(-9,7) and (-2,-8)
Substituting into:
m = (y2-y1)/(x2-x1)
m = (-8-7)/(-2-(-9))
m = (-15)/(-2+9)
m = (-15)/(7)
m = -15/7